From: STOSC::SMM "On-Duty SMM x5074" 5-JAN-1993 11:44:19.39 To: @[.DIS]HST_STAT CC: Subj: HST Status Report, 01/04 - 01/05. HUBBLE SPACE TELESCOPE DAILY REPORT #796 PERIOD COVERED: 0800 EST 01/04/93 - 0800 EST 01/05/93 Daily Status Report as of 005/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed S/C 4783 (DF-224 Version 8.3 Flight Software Installation) FSW 8.3 was successfully installed and is currently under going regression testing. FSW 8.3 contains the fine lock recentering algorithm which was designed to reduce fine lock loss of lock problems. No problems have occurred and the recentering algorithm is working as expected. 1.2 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.3 Completed HSP/UV2/POL 3234 (X-Ray Binaries) The High Speed Photometer was used to obtain eleven photometric and polarmetric light curve observations of the x-ray binary star HD77581. All eleven observations were examined in the OSS. The onboard target acquisitions were successful. There was a loss of lock which affected the last two observations. Otherwise, the observations looked good, and had a strong signal. 1.4 Completed Two Sets of WFC 3172 (Internal Monitor Proposal) The Wide Field Camera was used to take a series of internal dark, and bias observations. There were no external targets. The observations were executed as scheduled, and no problems were reported. 1.5 Completed WFC 2227 (Determine. of the Extragalactic Distance Scale) The Wide Field Camera was used to take a series of images of galaxy NGC-3031 (aka, M-81, a large spiral galaxy in Ursa Major) looking for Cepheid variable stars. This is to help improve our knowledge of the extragalactic distance scale. The observations were executed as scheduled, and no problems were reported. 1.6 Completed PC 4093 (A Snapshot of the Nuclear Regions) The Planetary Camera was used to take high resolution images of the inner regions of Markarian galaxy MARK74. The OSS monitored the observations, and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 4 Successful : 4 Losses of Lock : 1 (LOL at EOD 005/0514Z) 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 3605 (Stellar Winds and Post Main Sequence Evolution) 3.2 Complete Two Sets of WFC-PC 3172 (Internal Monitor Proposal) 3.3 Complete Two Sets of HSP/UV1 1091 (Ultraviolet Pulsations from X-Ray Pulsars) 3.4 Complete PC 4084 (Globular Cluster Cores, GTO Augmentation) 3.5 Complete FGS 4200 (Test of Galactic Escape Velocity at R(sun)) 3.6 Uplink Updated Ephemeris Tables, Per 2871 3.7 Complete WFC 3172 (Internal Monitor Proposal) 3.8 Complete HSP/UV2/POL 3234 (X-Ray Binaries) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time -END-